OpenStudy (anonymous):

A restaurant charges $8 for a large cheese pizza plus $1.50 per topping. Another restaurant charges $11 for a large cheese pizza plus $0.75 per topping. To find out how many toppings it would take for the pizzas to cost the same at both restaurants, which equation would you set up?

OpenStudy (anonymous):

8 + 1.5x = 11 + 0.75 x 8 + 0.75x = 11 + 1.5x 8x + 1.5 = 0.75x + 11 8x + 1.5 = 11x + 0.75

OpenStudy (anonymous):

8 + 1.5x = 11 + 0.75 x is what I got?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

& then it asks... For the same pizza restaurants referred to in the last question, how many toppings does a pizza have that costs the same at both restaurants?

OpenStudy (anonymous):

just solve that for x

OpenStudy (anonymous):

then that will find the amount of toppings that the pizza will have for it to be the same price

OpenStudy (anonymous):

I got 4 :o

OpenStudy (anonymous):

me too :)

OpenStudy (anonymous):

thanks again {:

OpenStudy (anonymous):

np :)